OpenStudy (anonymous):

ct all combinations of the following numbers that when added together would be rational. −6, 32√, 23, 13, 7√ Question 6 options: −6 and 23 32√ and 23 23 and 13 13 and 7√

OpenStudy (anonymous):

NO GUESS OR I THINK I NEED A WALK THROUGH OR DIRECT ANSWER

OpenStudy (kayders1997):

It's hard to read what is going on with the square roots and everything, can you rewrite it?

OpenStudy (anonymous):

oh ya sure

OpenStudy (anonymous):

\[its \sqrt{32} and \sqrt{7}\]

OpenStudy (kayders1997):

okay

OpenStudy (kayders1997):

first of all do you know what a rational number is

OpenStudy (anonymous):

oh shoot yea i know a rational number is and i need to fix the choices hang on

OpenStudy (anonymous):

\[-6, 2\sqrt{3}, 2/3, 1/3, and \sqrt{7}\]

OpenStudy (anonymous):

so the 23 and the 13 are fractions and 32 is 2squareroot3

OpenStudy (kayders1997):

okay

OpenStudy (kayders1997):

so those are the answer choice so it says you need to combine so you need to add those values together, what do you get when you add the first choices

OpenStudy (kayders1997):

so -6+2/3 I think your telling me based on the changes you said

OpenStudy (anonymous):

yes and it is -16/3 or -5 1/3

OpenStudy (kayders1997):

Yes does that have a end point (integer), or a repeating decimal?

OpenStudy (kayders1997):

I should have said whole number my bad not end point

OpenStudy (kayders1997):

do you understand what I am saying?

OpenStudy (anonymous):

so either integer or repeating decimal

OpenStudy (kayders1997):

Yes

OpenStudy (anonymous):

the decimal is repeating 5.3 reapeating

OpenStudy (kayders1997):

yes it is so that would be considered a combination that is rational

OpenStudy (kayders1997):

There is one more here

OpenStudy (anonymous):

ok

OpenStudy (kayders1997):

Do you think the square roots will give you an integer or repeating decimal?

OpenStudy (anonymous):

Yes?

OpenStudy (kayders1997):

most likely not unless you have a perfect square

OpenStudy (kayders1997):

Think about the 3rd function if you are adding 2/3+1/3 what do you get

OpenStudy (anonymous):

1

OpenStudy (kayders1997):

Yes is that an integer?

OpenStudy (anonymous):

yes

OpenStudy (kayders1997):

Yep so those are your two answers :)

OpenStudy (anonymous):

hey kayder can u help on my word problem @kayders1997